OpenStudy (anonymous):

Find the 12th partial sum of the summation of negative 7i plus 22, from i equals 1 to infinity.

OpenStudy (anonymous):

.25 40 -348 72

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):



OpenStudy (anonymous):

@vzfreakz

OpenStudy (yanasidlinskiy):

\(\sum_{i=1}^{12} (7i+22)=7\sum_{i=1}^{12}i+\sum_{i=1}^{12}22=7\left[\frac{12(13)}{2}\right]+12(22)\) <----I think that's what it would be.

OpenStudy (anonymous):

Whao Whoa. ;-; Whoa. Wait. What. @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

You've never seen this before?

OpenStudy (anonymous):

Uh, No ;-;

OpenStudy (yanasidlinskiy):

This is algebra. Correct?

OpenStudy (anonymous):

Yeah, Algebra 2 xD

OpenStudy (yanasidlinskiy):

Ok:) Wait..Does it look like this? \(\sum_{i=4}^{15} \) Btw, 15 should be on top and i-4 on bottom.

OpenStudy (yanasidlinskiy):

To actually calculate it, \(\sum_{i=4}^{15}( -2i-10)=-2 \sum_{i=4}^{15}i-\sum_{i=4}^{15}10\)

OpenStudy (anonymous):

It looks like the drawing up there xD if yah scroll up ;3

OpenStudy (yanasidlinskiy):

ok, yea. I didn't really notice...but anyways...This is what it is:) \(\sum_{i=4}^{15} (-2i-10)=-348\)

OpenStudy (anonymous):

Oh thanks GURL :3 Can you help with a few more? ;o

OpenStudy (yanasidlinskiy):

Close this and open a new one.